OpenStudy (anonymous):

What is the simplified form of the quantity 2 x minus 8 all over x minus 4 ? HELP

OpenStudy (anonymous):

(2x - 8)/(x -4) you can common factor out the 2 in the numerator: (2x - 8) = 2(x -4) the new fraction: 2(x-4) / (x-4) (x - 4)s cancel: 2/1 =2

OpenStudy (anonymous):

hmm here are the optionns x - 4, with the restriction x does not equal 4 2, with the restriction x does not equal 4 x - 4, with the restriction x does not equal 2 4, with the restriction x does not equal 2

OpenStudy (anonymous):

it's the second, I should also mention that x does not equal 4 because 4-4=0 and that would make the denominator 0

OpenStudy (anonymous):

k thanks